OpenStudy (anonymous):

a regular hexagon has sides of 3 feet. What is the area of the hexagon

OpenStudy (anonymous):

\[A = \frac{ 3 *\sqrt{3} }{ 2 } s^2\] where s is the side length

OpenStudy (anonymous):

yours is better^

OpenStudy (anonymous):

still don't get it I am getting an error

OpenStudy (anonymous):

All you need to do is plug in your side length and do the multiplication. What are you getting an error with?

OpenStudy (anonymous):

ok I got 9 sq3 ft is that right

OpenStudy (anonymous):

9 sq root 3

OpenStudy (anonymous):

No. It isn't. written differntly it can look like this if it helps. \[s^2*\frac{ 3 }{ 2 } *\sqrt{3}\]

OpenStudy (anonymous):

ok I believe I got it 27root3

OpenStudy (anonymous):

Almost, do it part by part. s^2 = what and then multiple that by 3/2 = 1.5 then by root 3

OpenStudy (anonymous):

13.5

OpenStudy (anonymous):

13.5 *sqrt(3). Yes.

OpenStudy (anonymous):

yes that's what I meant thank you

OpenStudy (anonymous):

No problem. Glad to help.